OpenStudy (anonymous):

Todd bought 1/2 pound of chicken. He used of the chicken for sandwiches. How much chicken did Todd use for sandwiches? A. 1/6lb B. 1/3lb C. 2/5lb D. 2/3lb

OpenStudy (anonymous):

please help

OpenStudy (welshfella):

you missed out part of the quaestion:- 'He used ? of the checken'

OpenStudy (anonymous):

what fraction of the chicken did he use

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

he used 2/3

OpenStudy (mathmale):

Please type in the whole question again; cut and paste as necessary.

OpenStudy (anonymous):

1/3 is answer

OpenStudy (welshfella):

Oh so its 2/3 of 1/2 lb = 2/3 * 1/2 lb

OpenStudy (anonymous):

He used x part of chicken in his sandwich so how much chicken he used b = a.x where a = 1/2 pound. Can you find it and understand whats happening.

OpenStudy (welshfella):

@QUestionerperson plz dont give a direct answer

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

thanks guys you all earned a fan :P